OpenStudy (anonymous):

v = 1/3 nr^2 h find the value of h if n = 3, r = 4, v = 48

OpenStudy (anonymous):

okay so the problem would look like 48=1/3 * 3 * 4^2 *h

OpenStudy (anonymous):

now just follow order of operations

OpenStudy (anonymous):

so first we do the power? r^2 = 16 16 x n (which is 3) = 48 This is what I am getting confused on, it equals 48, whereas V = 48 :/

OpenStudy (anonymous):

@Rushwr

OpenStudy (petiteme):

\[48 = \frac{ 1 }{ 3 } * 3 * 4^{2}* h\]

OpenStudy (petiteme):

You said to do the power first, right? So, it'll be... \[48=\frac{ 1 }{ 3 }*3*16*h\]

OpenStudy (petiteme):

Perform the operation :)

OpenStudy (anonymous):

@petiteme yes that's right, however i need to find the value of h. it's proving to be confusing

OpenStudy (petiteme):

Okay. Do it step by step. Multiply 1/3 to 3 first.

OpenStudy (anonymous):

16 x 3 = 48. one third of 48 = 16

OpenStudy (petiteme):

Okay that's good. So it'll be. 48 = 16h

OpenStudy (anonymous):

Are you sure? LOOL that seems so easy I feel dumb now, here's me trying to dvide it again

OpenStudy (petiteme):

Yeah lol.

OpenStudy (petiteme):

You didn't get the value of h yet

OpenStudy (anonymous):

@petiteme LOOL OH yes

OpenStudy (petiteme):

Do you know how to get it?

OpenStudy (anonymous):

no, is it 16 didivded by 3

OpenStudy (petiteme):

No. So our equation now is 48 = 16h Divide both sides by 1/16, so that it'll be cancelled and the h remain. Hope that makes sense huhu

OpenStudy (anonymous):

I got it

OpenStudy (anonymous):

Its 3

OpenStudy (petiteme):